OpenStudy (anonymous):

Could anybody help me solve this question? What is the solution set of 2x^2 + x = 15? {3, -5/2} {-3, 5/2} {2, -3/2}

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Could you show me how you got that?

OpenStudy (anonymous):

i can

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

take 15 in left side 2x^2 -x -15 = 0

OpenStudy (anonymous):

then 2x^2 -6x + 5x - 15 = 0

OpenStudy (anonymous):

take common 2x (x -3) +5 ( x - 3 )

OpenStudy (anonymous):

(2x + 5 ) ( x - 3 )

OpenStudy (anonymous):

2x + 5 = 0 2x = -5 x = -5 / 2

OpenStudy (anonymous):

x-3 = 0 x=3

OpenStudy (anonymous):

did this help or not